OpenStudy (anonymous):

Using Calculus techniques, find the approximate value of:

OpenStudy (anonymous):

\[\sqrt[3]{25}\]

OpenStudy (amistre64):

find an approximate cuberoot; and keep reiterating? I know there is a certain "way" you can get to it.... but I forget it right off hand.

OpenStudy (amistre64):

f(x) = f'(x) + ....

OpenStudy (amistre64):

x^(1/3) gets us f'(x) = (1/3)x(-2/3) 1 --------- = the slope of your "line equation" cbrt(x^2)

OpenStudy (amistre64):

y = 1/cbrt(x^2) + b i think is the way it can go...then find out what this value is for x and reiterate it again tilll you get to a level of accuracy that you are comfortable with